OpenStudy (purple_pink):

50 is what percent of 40?

OpenStudy (purple_pink):

@johnweldon1993

OpenStudy (keepcalmandqson):

110%

OpenStudy (emmanuel77m):

No its 125%

OpenStudy (purple_pink):

thank you! :)

OpenStudy (purple_pink):

which one?

OpenStudy (emmanuel77m):

125%

OpenStudy (purple_pink):

@rebeccaxhawaii

OpenStudy (emmanuel77m):

50/40 X 100 = 5000/40 = 500/4 = 125% ANSWER

OpenStudy (fortytherapper):

Right, 40 would be 100% in this case. So setting it up \[\frac{ 50 }{ 40 }=\frac{ x }{ 100 }\]

OpenStudy (anonymous):

50 - 40 = 10. Then 10/40 = 25% So if 40 is 100%, then 100%+25% = 125%

OpenStudy (keepcalmandqson):

sorry i got messed up in my question

OpenStudy (purple_pink):

it's ok! :) we all make mistakes *thumbs up*